OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12

OpenStudy (anonymous):

A. y = -12 - 4/3 x B. y = -4 - 4/3 x C. y = -12 + 4x D. y = -4 + 4x

OpenStudy (anonymous):

What could you do to get y isolated in one side?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

What could you do to move 4x to the right side?